Solve the equation by graphing. x-y=4 2x-5y=2
Answer by mananth(16949)

Find the x & y intercept points by plugging x=0,y=0 in both equations.
You can also take any x value which will give y values which are easily plotted
equation I
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= 4 ( 4 , 0 )
Equation 2
when x= 0 y= -0.4 ( 0 , -0.4 )
when y = 0 x= 1 ( 1 , 0 )



You will observe that the point of intersection is 6 , 2
The solution of the system is ( 6 , 2 )
m.ananth@hotmail.ca